# Physics-informed machine learning for combustion: A review

### v2

@@ 7.1 Datasets and benchmarks
It is known that data, algorithms, and computing power are the three driving forces behind AI development, reflecting the important role of data in AI. Specifically, data are the foundation for training models and their quality and quantity can directly influence the model performance. Therefore, it is important to establish public datasets and related benchmarks which can reduce redundant work, facilitate reproducibility, and standardize model evaluation and comparison. By definition, a benchmark is composed of the dataset(s) on which the benchmark is trained or evaluated, and a reference implementation in any language 354. Note that although PIML for forward problems is purely physics-driven and data-free, ground-truth data is also necessary for model evaluation, especially in academia. A large number of datasets and benchmarks have been established for various AI tasks, with the most famous of which is the ImageNet 355 for CV. In contrast, the datasets and benchmarks for SciML and fluid, and combustion are fewer, but have been growing rapidly in recent years. Compared to traditional AI datasets, datasets for SciML, fluid, and combustion usually exhibit unique properties due to their physics-driven nature, such as high-dimensional, multiscale, costly-generated, and interdisciplinary, which pose challenges for the generation, storage, sharing, and use of the datasets.
@@
Table 6 lists some representative public datasets and benchmarks for SciML, fluid, and combustion. Since fluid mechanics is an important topic of SciML, fluid datasets are included in many SciML dataset/benchmarks, such as those listed in Table 6. Takamoto et al. 141 developed the PDEBench, providing 35 datasets based on 11 well-known time-dependent and time-independent PDEs and evaluating popular SciML models, such as the FNO, U-Net, and PINNs, in an AR way (see Section 2.3.1). Luo et al. 139 constructed the CFDBench for evaluating the NOs on four CFD cases in both AR and NAR ways. Koehler et al. 140 established the APEBench not only for evaluating AR neural emulators, bust also providing an integrated differentiable simulation framework. As a PINN-specific benchmark, PINNacle developed by Hao et al. 356 provides a diverse dataset covering over 20 distinct PDEs from various domains. PINNacle also serves as a user-friendly toolbox that incorporate more than ten PINN methods, some of which are introduced in Section 2.3. A detailed comparison of fluid-contained SciML dataset/benchmarks can be found in the paper of FlowBench 357. Note that although some SciML benchmarks evaluate the models on AR tasks, their datasets can also be used for evaluating the classical NAR PIML models. In addition to the fluid data in SciML dataset/benchmarks, there are also some classical fluid datasets in the fluid community, such as the well-known Johns Hopkins turbulence database (JHTDB) 358 and the ERCOFTAC (European Research Community On Flow, Turbulence, And Combustion) classic collection database 359. They were originally established for use in fluid tasks, such as turbulence modeling, but now they can also be used for SciML. A list of fluid databases and collections from the fluid community can be found in Reference 360. In addition to datasets and benchmarks, knowledge bases also serve as important resources for both academia and industry. An example is the ERCOFTAC knowledge base wiki 361, which provides structured knowledge and advice for CFD, covering non-reacting and reacting flows.
@@
ScalarFlow 362 A large-scale volumetric dataset of real-world scalar transport flows URL
@@
AirfRANS 363 A dataset of solutions of 2D incompressible RANS equations over airfoils URL
PINNacle 356 A comprehensive benchmark of PINNs for solving PDEs URL
@@
FlowBench 357 A large-scale benchmark for flow simulation over complex geometries URL
@@
The Well 364 A large-scale simulation dataset of diverse spatiotemporal physical systems URL
JHTDB 358 Johns Hopkins turbulence database URL
ERCOFTAC 359 ERCOFTAC classic collection database URL
NASA TMR 365 Turbulence modeling resource of NASA Langley Research Center URL
Towne et al. 366 A database for reduced-complexity modeling of fluid flows URL
TNF library 367 A library of well-documented flames established by the TNF workshop URL
BLASTNet 368, 369, 336 A large-scale community-involved network-of-datasets of reacting and non-reacting flows obtained from DNS for benchmarking ML methods URL
GlobFire 370 A global wildfire dataset for the analysis of fire regimes and fire behaviour URL
Sim2Real-Fire 371 A multi-modal simulation dataset for forecast and backtracking of real-world forest fire URL
FireBench 372 An ensemble of LES dataset of 3D wildfire spread URL
For combustion, the datasets and benchmarks are fewer than those for SciML and fluid. The TNF workshop is an international collaboration on turbulent combustion and established a library of well-documented flames that includes multiscalar and velocity data 367. The flame types cover from simple jet flames to piloted jet flames, bluff body flames, and swirl flames. Some classic flames such as the Sandia flames 329 are also included and have already been studied using PIML 311, 312 (see Section 5.3). Chung et al. 368, 369 established the BLASTNet (Bearable Large Accessible Scientific Training Network-of-datasets), a network-of-datasets of reacting and non-reacting flows which combines community involvement, public data repositories, and lossy compression algorithms. Figure 31 gives an overview of the BLASTNet, which is still developing. Using the DNS data of a H2-air lifted flame, the authors showcased the usage of BLASTNet by benchmarking the CNNs on both the tasks of combustion regime classification and filtered reaction rate regression 369. Later, the BLASTNet 2.0 336 was established with extended samples, configurations, and size, and has been used for benchmarking the scaling behaviour of five DL models on the 3D volumetric SR task. One of the five models adopted the physics-informed loss in a broad sense, as mentioned in Section 5.5. BLASTNet has also been used for other PIML tasks, such as the PiMAPNet 119 introduced in Section 5.3. Table 6 also provides some datasets of fire, including the GlobFire 370, Sim2Real-Fire 371, and FireBench 372. The FireBench has also been uploaded to the BLASTNet. For a detailed comparison of datasets for wildfire analysis, readers can refer to the paper of Sim2Real-Fire 371.
Figure 31: Overview of the BLASTNet 368, 369, 336.
@@ 7.2 Simulation tools (data generators)
In SciML, even with the growing number of public datasets and benchmarks, researchers still need to generate their own datasets using simulation tools, which are essential for customizing domain-specific cases, controlling data quality, and filling the inherent sparsity of public datasets. Table 7 gives a non-exhaustive list of public tools for ODE/PDE solution, fluid simulation, and combustion simulation. There are no clear boundaries between these three types of tools: Fluid simulation often falls under PDE solution, while combustion simulation often falls under either ODE solution or fluid simulation. Some ODE/PDE solution tools can be used for fluid/combustion simulation, while others cannot. For example, The CVODE 373 in SUNDIALS 374, 375 has been used in many aforementioned combustion researches 292, 279, 117, 118. The PyClaw 376, Python version of Clawpack 377, 378, has been used for simulating supersonic combustion 314. For fluid simulation, the OpenFOAM 379, 380, 381, 382 is a widely-used open-source CFD software and has many solvers for combustion, such as the chemFoam, reactingFoam, fireFoam, flameletFoam 383, and FGMFoam 384. Recently, researchers have developed tools that utilize the large language models (LLMs) to automate OpenFOAM programming, including the MetaOpenFOAM 385, 386 and OpenFOAMGPT 387, 388. Cantera 389 is an open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes, and has been used in PIML for 1D laminar flames 115. Some AI-integrated simulators are also provided in Table 7, including (1) the DeepFlame 118, a deep learning empowered open-source platform for reacting flow simulations, which coupled the three tools of OpenFOAM, Torch 390, and Cantera, and (2) differentiable solvers, including the torchdiffeq 391, $\bf{\Phi}_{\textrm{Flow}}$ 392, 393, JAX-Fluids 394, 395, Arrhenius.jl 396, and JANC 397. Table 7 also lists some tools that can be used for or are specifically designed for the DNS of fluid and combustion, including the Dedalus 398, Nek5000 399, SENGA+ 400, and AVBP 401. Note that most DNS tools for combustion are in-house codes, such as the NGA 402, S3D 403, and Athena-RFX 404, which have been used for establishing the BLASTNet 336. See more examples in this review 405 and the paper of BLASTNet 2.0 336. However, considering the extremely high cost of DNS for combustion, it is recommended to directly use the public datasets mentioned in Section 7.1 rather than generating them, especially for AI-community researchers.
@@
The tools in Table 7 have been used for generating datasets and benchmarks in Table 6. For example, the Clawpack and $\bf{\Phi}_{\textrm{Flow}}$ were used to establish the PDEBench 141, while the Clawpack and Dedalus were used to establish the Well 364. In addition to free tools, researchers can also use some widely-used commercial software for fluid/combustion simulation and dataset generation, such as the ANSYS Fluent 406, Chemkin 407, and COMSOL Multiphysics 408. For example, COMSOL Multiphysics has been used for generating data in PIML for combustion 342 and generating the datasets of PINNacle 356. An overview of both free and commercial CFD software can be found in Reference 409.
@@
SciPy 410 An open-source scientific computing library in Python language URL
SUNDIALS 374, 375 An open-source SUite of Nonlinear and DIfferential/ALgebraic Equation Solvers URL
GEKKO 411 A Python package for ML and optimization of mixed-integer and DAEs URL
@@
FEniCS 412, 413, 414 An open-source computing platform for solving PDEs based on the FEM URL
FreeFEM++ 415 A free software for solving PDEs based on the FEM URL
FiPy 416 A PDE solver using Python based on the FVM URL
Clawpack 377, 378, 376 A package that collects FVMs for hyperbolic systems of conservation laws URL
Dedalus 398 An open-source framework for solving PDEs using spectral methods URL
OpenFOAM 379, 380, 381 An open-source CFD software URL
Nek5000 399 A fast and scalable open-source CFD solver using the spectral element method URL
$\bf{\Phi}_{\textrm{Flow}}$ 392, 393 A differentiable PDE solving framework for optimization and ML applications URL
JAX-Fluids 394, 395 A differentiable CFD solver for 3D compressible single-phase and two-phase flows URL
CFD Python 417 An online learning module for CFD with Python URL
Cantera 389 An open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes URL
FlameMaster 418 An open source C++ package for 0D and 1D laminar combustion calculations URL
OpenSMOKE++ 419 A framework for simulating reacting systems with detailed kinetic mechanisms URL
SENGA+ 400 A DNS code for 3D compressible turbulent reacting flows URL
AVBP 401 A suite of CFD tools for performing LES/DNS of compressible reacting flows URL
@@
Arrhenius.jl 396 A differentiable combustion simulation package URL
JANC 397 A differentiable compressible reacting flow solver URL
FDS 420 Fire dynamics simulator, a LES code for fire-driven fluid flows URL
@@ 7.3 PIML tools
There are many tools for implementing the classical PIML, i.e., PINNs, some of which are presented in Table 8. Most tools are written in Python, except for the ADCME 421 and NeuralPDE 422, which are written in Julia. Most tools can serve as solvers in which users only need to define the problem, while other tools, such as the SciANN 423 and ADCME 421, work as wrappers in which users still need to deal with most of the implementation details. Since AD has been implemented in modern AI frameworks, such as the PyTorch 424, 425, TensorFlow 426, and Keras 427, they are usually used as the backend for these PIML tools. Most PIML tools support one or two backends, except for the DeepXDE 170, which supports the PyTorch, TensorFlow, JAX 428, and PaddlePaddle 429. For more details of these PIML tools, readers can refer to these reviews 68, 80, 94, 92, where Reference 92 also listed the merits and limitations of some tools.
@@
PyDEns 430 A Python framework for solving ODE/PDEs using NNs URL
NeuroDiffEq 431, 432 A Python package for solving DEs with NNs URL
NVIDIA PhysicsNeMo 433 An open-source framework for building, training, and fine-tuning SciML models URL
Elvet 434 An NN-based DE and variational problem solver URL
@@
IDRLnet 436 A PINN library on top of PyTorch URL
PINA 437 A Python library designed to simplify and accelerate the development of SciML URL
SciANN 423 A Keras/TensorFlow wrapper for scientific computations and PIML using NNs URL
ADCME 421 A Julia framework to solve inverse problems involving physical simulations and NNs URL
NeuralPDE 422 A Julia package for solving PDEs using PINNs URL
Some PIML tools have been used in PIML for combustion. For example, the NVIDIA PhysicsNeMo 431, previously known as NVIDIA Modulus and SimNet 438, was used in Reference 307. The DeepXDE was used for implementing the FlamePINN-1D 115 and optimizing the extended VDP system 346. Some studies also implement the PIML for combustion directly using the modern AI frameworks, including the PyTorch 310, 308, 309, 314 and TensorFlow 312, which is more flexible but cumbersome than using PIML tools. To show how popular PIML tools can be used for combustion research, a well-commented code demo is provided in Listing 32, which is based on the DeepXDE and adapted from the codes of FlamePINN-1D. Taking the 2D combustion reacting flow as example, the demo shows how to code the Equations (3.38) to (3.42) using a PIML tool, in which the convenience of AD to compute derivatives is clearly demonstrated. To illustrate how to infer unknown parameters in inverse problems, the demo also defines an unknown correction coefficient for viscosity. It is straightforward to modify the demo for 3D combustion. For more details about the implementation of IC/BC/OCs, network, training, and evaluation, and how to code other equations in Section 3 in a differentiable way, readers can refer to the codes of FlamePINN-1D 115.

### v3

@@ 7.1 Datasets and benchmarks
It is known that data, algorithms, and computing power are the three driving forces behind AI development, reflecting the important role of data in AI. Specifically, data are the foundation for training models and their quality and quantity can directly influence the model performance. Therefore, it is important to establish public datasets and related benchmarks which can reduce redundant work, facilitate reproducibility, and standardize model evaluation and comparison. By definition, a benchmark is composed of the dataset(s) on which the benchmark is trained or evaluated, and a reference implementation in any language 356. Note that although PIML for forward problems is purely physics-driven and data-free, ground-truth data is also necessary for model evaluation, especially in academia. A large number of datasets and benchmarks have been established for various AI tasks, with the most famous of which is the ImageNet 357 for CV. In contrast, the datasets and benchmarks for SciML and fluid, and combustion are fewer, but have been growing rapidly in recent years. Compared to traditional AI datasets, datasets for SciML, fluid, and combustion usually exhibit unique properties due to their physics-driven nature, such as high-dimensional, multiscale, costly-generated, and interdisciplinary, which pose challenges for the generation, storage, sharing, and use of the datasets.
@@
Table 6 lists some representative public datasets and benchmarks for SciML, fluid, and combustion. Since fluid mechanics is an important topic of SciML, fluid datasets are included in many SciML dataset/benchmarks, such as those listed in Table 6. Takamoto et al. 141 developed the PDEBench, providing 35 datasets based on 11 well-known time-dependent and time-independent PDEs and evaluating popular SciML models, such as the FNO, U-Net, and PINNs, in an AR way (see Section 2.3.1). Luo et al. 139 constructed the CFDBench for evaluating the NOs on four CFD cases in both AR and NAR ways. Koehler et al. 140 established the APEBench not only for evaluating AR neural emulators, bust also providing an integrated differentiable simulation framework. As a PINN-specific benchmark, PINNacle developed by Hao et al. 358 provides a diverse dataset covering over 20 distinct PDEs from various domains. PINNacle also serves as a user-friendly toolbox that incorporate more than ten PINN methods, some of which are introduced in Section 2.3. A detailed comparison of fluid-contained SciML dataset/benchmarks can be found in the paper of FlowBench 359. Note that although some SciML benchmarks evaluate the models on AR tasks, their datasets can also be used for evaluating the classical NAR PIML models. In addition to the fluid data in SciML dataset/benchmarks, there are also some classical fluid datasets in the fluid community, such as the well-known Johns Hopkins turbulence database (JHTDB) 360 and the ERCOFTAC (European Research Community On Flow, Turbulence, And Combustion) classic collection database 361. They were originally established for use in fluid tasks, such as turbulence modeling, but now they can also be used for SciML. A list of fluid databases and collections from the fluid community can be found in Reference 362. In addition to datasets and benchmarks, knowledge bases also serve as important resources for both academia and industry. An example is the ERCOFTAC knowledge base wiki 363, which provides structured knowledge and advice for CFD, covering non-reacting and reacting flows.
@@
ScalarFlow 364 A large-scale volumetric dataset of real-world scalar transport flows URL
@@
AirfRANS 365 A dataset of solutions of 2D incompressible RANS equations over airfoils URL
PINNacle 358 A comprehensive benchmark of PINNs for solving PDEs URL
@@
FlowBench 359 A large-scale benchmark for flow simulation over complex geometries URL
@@
The Well 366 A large-scale simulation dataset of diverse spatiotemporal physical systems URL
JHTDB 360 Johns Hopkins turbulence database URL
ERCOFTAC 361 ERCOFTAC classic collection database URL
NASA TMR 367 Turbulence modeling resource of NASA Langley Research Center URL
Towne et al. 368 A database for reduced-complexity modeling of fluid flows URL
TNF library 369 A library of well-documented flames established by the TNF workshop URL
BLASTNet 370, 371, 334 A large-scale community-involved network-of-datasets of reacting and non-reacting flows obtained from DNS for benchmarking ML methods URL
GlobFire 372 A global wildfire dataset for the analysis of fire regimes and fire behaviour URL
Sim2Real-Fire 373 A multi-modal simulation dataset for forecast and backtracking of real-world forest fire URL
FireBench 374 An ensemble of LES dataset of 3D wildfire spread URL
For combustion, the datasets and benchmarks are fewer than those for SciML and fluid. The TNF workshop is an international collaboration on turbulent combustion and established a library of well-documented flames that includes multiscalar and velocity data 369. The flame types cover from simple jet flames to piloted jet flames, bluff body flames, and swirl flames. Some classic flames such as the Sandia flames 329 are also included and have already been studied using PIML 309, 311 (see Section 5.3). Chung et al. 370, 371 established the BLASTNet (Bearable Large Accessible Scientific Training Network-of-datasets), a network-of-datasets of reacting and non-reacting flows which combines community involvement, public data repositories, and lossy compression algorithms. Using the DNS data of a H2-air lifted flame, the authors showcased the usage of BLASTNet by benchmarking the CNNs on both the tasks of combustion regime classification and filtered reaction rate regression 371. Later, the BLASTNet 2.0 334 was established with extended samples, configurations, and size, and has been used for benchmarking the scaling behaviour of five DL models on the 3D volumetric SR task. One of the five models adopted the physics-informed loss in a broad sense, as mentioned in Section 5.5. BLASTNet has also been used for other PIML tasks, such as the PiMAPNet 119 introduced in Section 5.3. Table 6 also provides some datasets of fire, including the GlobFire 372, Sim2Real-Fire 373, and FireBench 374. The FireBench has also been uploaded to the BLASTNet. For a detailed comparison of datasets for wildfire analysis, readers can refer to the paper of Sim2Real-Fire 373.
@@ 7.2 Simulation tools (data generators)
In SciML, even with the growing number of public datasets and benchmarks, researchers still need to generate their own datasets using simulation tools, which are essential for customizing domain-specific cases, controlling data quality, and filling the inherent sparsity of public datasets. Table 7 gives a non-exhaustive list of public tools for ODE/PDE solution, fluid simulation, and combustion simulation. There are no clear boundaries between these three types of tools: Fluid simulation often falls under PDE solution, while combustion simulation often falls under either ODE solution or fluid simulation. Some ODE/PDE solution tools can be used for fluid/combustion simulation, while others cannot. For example, The CVODE 375 in SUNDIALS 376, 377 has been used in many aforementioned combustion researches 292, 279, 117, 118. The PyClaw 378, Python version of Clawpack 379, 380, has been used for simulating supersonic combustion 313. For fluid simulation, the OpenFOAM 381, 382, 383, 384 is a widely-used open-source CFD software and has many solvers for combustion, such as the chemFoam, reactingFoam, fireFoam, flameletFoam 385, and FGMFoam 386. Recently, researchers have developed tools that utilize the large language models (LLMs) to automate OpenFOAM programming, including the MetaOpenFOAM 387, 388 and OpenFOAMGPT 389, 390. Cantera 391 is an open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes, and has been used in PIML for 1D laminar flames 115. Some AI-integrated simulators are also provided in Table 7, including (1) the DeepFlame 118, a deep learning empowered open-source platform for reacting flow simulations, which coupled the three tools of OpenFOAM, Torch 392, and Cantera, and (2) differentiable solvers, including the torchdiffeq 393, $\bf{\Phi}_{\textrm{Flow}}$ 394, 395, JAX-Fluids 396, 397, Arrhenius.jl 398, and JANC 399. Table 7 also lists some tools that can be used for or are specifically designed for the DNS of fluid and combustion, including the Dedalus 400, Nek5000 401, SENGA+ 402, and AVBP 403. Note that most DNS tools for combustion are in-house codes, such as the NGA 404, S3D 405, and Athena-RFX 406, which have been used for establishing the BLASTNet 334. See more examples in this review 407 and the paper of BLASTNet 2.0 334. However, considering the extremely high cost of DNS for combustion, it is recommended to directly use the public datasets mentioned in Section 7.1 rather than generating them, especially for AI-community researchers.
@@
The tools in Table 7 have been used for generating datasets and benchmarks in Table 6. For example, the Clawpack and $\bf{\Phi}_{\textrm{Flow}}$ were used to establish the PDEBench 141, while the Clawpack and Dedalus were used to establish the Well 366. In addition to free tools, researchers can also use some widely-used commercial software for fluid/combustion simulation and dataset generation, such as the ANSYS Fluent 408, Chemkin 409, and COMSOL Multiphysics 410. For example, COMSOL Multiphysics has been used for generating data in PIML for combustion 339 and generating the datasets of PINNacle 358. An overview of both free and commercial CFD software can be found in Reference 411.
@@
SciPy 412 An open-source scientific computing library in Python language URL
SUNDIALS 376, 377 An open-source SUite of Nonlinear and DIfferential/ALgebraic Equation Solvers URL
GEKKO 413 A Python package for ML and optimization of mixed-integer and DAEs URL
@@
FEniCS 414, 415, 416 An open-source computing platform for solving PDEs based on the FEM URL
FreeFEM++ 417 A free software for solving PDEs based on the FEM URL
FiPy 418 A PDE solver using Python based on the FVM URL
Clawpack 379, 380, 378 A package that collects FVMs for hyperbolic systems of conservation laws URL
Dedalus 400 An open-source framework for solving PDEs using spectral methods URL
OpenFOAM 381, 382, 383 An open-source CFD software URL
Nek5000 401 A fast and scalable open-source CFD solver using the spectral element method URL
$\bf{\Phi}_{\textrm{Flow}}$ 394, 395 A differentiable PDE solving framework for optimization and ML applications URL
JAX-Fluids 396, 397 A differentiable CFD solver for 3D compressible single-phase and two-phase flows URL
CFD Python 419 An online learning module for CFD with Python URL
Cantera 391 An open-source suite of tools for problems involving chemical kinetics, thermodynamics, and transport processes URL
FlameMaster 420 An open source C++ package for 0D and 1D laminar combustion calculations URL
OpenSMOKE++ 421 A framework for simulating reacting systems with detailed kinetic mechanisms URL
SENGA+ 402 A DNS code for 3D compressible turbulent reacting flows URL
AVBP 403 A suite of CFD tools for performing LES/DNS of compressible reacting flows URL
@@
Arrhenius.jl 398 A differentiable combustion simulation package URL
JANC 399 A differentiable compressible reacting flow solver URL
FDS 422 Fire dynamics simulator, a LES code for fire-driven fluid flows URL
@@ 7.3 PIML tools
There are many tools for implementing the classical PIML, i.e., PINNs, some of which are presented in Table 8. Most tools are written in Python, except for the ADCME 423 and NeuralPDE 424, which are written in Julia. Most tools can serve as solvers in which users only need to define the problem, while other tools, such as the SciANN 425 and ADCME 423, work as wrappers in which users still need to deal with most of the implementation details. Since AD has been implemented in modern AI frameworks, such as the PyTorch 426, 427, TensorFlow 428, and Keras 429, they are usually used as the backend for these PIML tools. Most PIML tools support one or two backends, except for the DeepXDE 170, which supports the PyTorch, TensorFlow, JAX 430, and PaddlePaddle 431. For more details of these PIML tools, readers can refer to these reviews 68, 80, 94, 92, where Reference 92 also listed the merits and limitations of some tools.
@@
PyDEns 432 A Python framework for solving ODE/PDEs using NNs URL
NeuroDiffEq 433, 434 A Python package for solving DEs with NNs URL
NVIDIA PhysicsNeMo 435 An open-source framework for building, training, and fine-tuning SciML models URL
Elvet 436 An NN-based DE and variational problem solver URL
@@
IDRLnet 438 A PINN library on top of PyTorch URL
PINA 439 A Python library designed to simplify and accelerate the development of SciML URL
SciANN 425 A Keras/TensorFlow wrapper for scientific computations and PIML using NNs URL
ADCME 423 A Julia framework to solve inverse problems involving physical simulations and NNs URL
NeuralPDE 424 A Julia package for solving PDEs using PINNs URL
Some PIML tools have been used in PIML for combustion. For example, the NVIDIA PhysicsNeMo 433, previously known as NVIDIA Modulus and SimNet 440, was used in Reference 304. The DeepXDE was used for implementing the FlamePINN-1D 115 and optimizing the extended VDP system 351. Some studies also implement the PIML for combustion directly using the modern AI frameworks, including the PyTorch 307, 305, 306, 313 and TensorFlow 311, which is more flexible but cumbersome than using PIML tools. To show how popular PIML tools can be used for combustion research, a well-commented code demo is provided in Listing 1, which is based on the DeepXDE and adapted from the codes of FlamePINN-1D. Taking the 2D combustion reacting flow as example, the demo shows how to code the Equations (3.38) to (3.42) using a PIML tool, in which the convenience of AD to compute derivatives is clearly demonstrated. To illustrate how to infer unknown parameters in inverse problems, the demo also defines an unknown correction coefficient for viscosity. It is straightforward to modify the demo for 3D combustion. For more details about the implementation of IC/BC/OCs, network, training, and evaluation, and how to code other equations in Section 3 in a differentiable way, readers can refer to the codes of FlamePINN-1D 115.